# A Generalized Back-Door Criterion for Linear Regression

### v4

@@ Abstract
What assumptions about the data-generating process are required to permit a causal interpretation of partial regression coefficients? To answer this question, this paper generalizes Pearl’s single-door and back-door criteria and proposes a new criterion that enables the identification of total or partial causal effects. In addition, this paper elucidates the mechanism of post-treatment bias, showing that a repeated sequence of nodes can be a potential source of this bias. The results apply to linear data-generating processes represented by directed acyclic graphs with distribution-free error terms.
@@ 1 Introduction
As a generalization of this approach, Pearl1995 developed a framework for causal inference that can also be applied to non-linear causal structures. Within this framework, Pearl1998 applied the back-door criterion—originally proposed by Pearl1993 as a method for covariate selection—to linear data-generating processes represented by directed acyclic graphs (DAGs), showing that it provides a sufficient condition under which a regression coefficient coincides with a total effect in the sense of SEM. The back-door criterion has since been generalized by Shpitser2012, Zander2014, Maathuis2015, Perkovic2015, and Perkovic2020. In addition to the back-door criterion, other approaches, such as the front-door criterion (Pearl1995) and conditional instrumental variables (BritoandPearl2002), have been established for identifying total effects.
Meanwhile, some previous studies have explored the identification of partial causal relationships. Pearl1998 proposed the single-door criterion for identifying direct effects, and Avin2005 and Shpitser2008 studied the identifiability of path-specific effects. BritoandPearl2002, Tian2009, Foygel2012, Chen2016, and Barber2022 discussed the identifiability of path coefficients by considering the paths that contribute to covariance. Fulcher2020 studied a generalized front-door criterion for identifying indirect effects. However, conditions guaranteeing that a partial regression coefficient coincides with a causal effect not mediated by covariates have been identified only in cases where the direct or total effect is of interest. Consequently, the logical validity of the causal interpretation of partial regression coefficients, as mentioned at the beginning, remains insufficiently established. Given that regression analysis is widely used not only in academic research but also in various practical contexts, it is crucial to provide a theoretically sound basis for its interpretation.
@@
  • Providing a graphical criterion under which a partial regression coefficient can be interpreted as a causal effect, controlling for the other explanatory variables at fixed values.
@@
The remainder of this paper is organized as follows. Section 2 introduces key terms and propositions as preliminaries. Section 3 formulates the problem addressed in this study and presents the main theorem. The section also discusses the mechanisms of post-treatment bias. Section 4 provides illustrative examples. All proofs are presented in Section 5.
@@ 2.4 Controlled Total Effects
Since independence among the error terms is not assumed, this equality cannot be derived based on do-calculus(Pearl2009). However, given that the notation involving the do-operator represents probabilistic concepts of the post-intervention distribution, this interpretation is appropriate. See Example 4.2 in the next section for a concrete illustration of this measure.
@@ 3.2 Main Results
These corollaries are originally established in Spirtes1998 and Pearl1998 based on conditional independence relationships and conditional covariances in linear Gaussian models. However, as also noted in Spirtes1998 and Nandy2017, their proofs are without loss of generality, since the partial regression coefficients depend solely on $(\bm{I}_{N}-\bm{A})^{-1}\mathop{\rm Var}\nolimits(\bm{u})(\bm{I}_{N}-\bm{A})^{-1}$. Here, $\bm{I}_{N}$ denotes the $N\times N$ identity matrix. A contribution of the present study is to show that linear causal effects not mediated by covariates can also be consistently estimated using ordinary least squares, even without assuming independence or normality of the error terms in the data-generating process.
@@ 3.3 Post-Treatment Bias
Biases that arise from conditioning on variables affected by a treatment (i.e., intervened variables) are referred to as post-treatment bias. Examining the mechanism of this bias helps clarify why condition (i) of the selective-door criterion is required.
@@
As indicated by the back-door criterion, confounding bias can be eliminated by blocking all back-door paths. The first condition (i) of the selective-door criterion implies that $\tau_{pj|do(\bm{S}_{2}\setminus\{X_{j}\})}=0$ and can be understood as a condition to prevent post-treatment bias. Furthermore, this perspective clarifies why merely blocking undirected paths is not sufficient to remove all sources of bias. Since a path is defined as a sequence of distinct nodes, for instance, the sequence of nodes $X_{j}\rightarrow X_{i}\rightarrow X_{p}\leftarrow X_{i}$ is not considered a path. However, the estimated effect of $X_{j}$ on $X_{i}$ can still be distorted through this structure if $X_{p}$ is conditioned on. Another instance of such a structure is provided in Example 4.1 below.

### v5

@@ Abstract
What assumptions about the data-generating process are required to permit a causal interpretation of partial regression coefficients? To answer this question, this paper generalizes Pearl’s single-door and back-door criteria and proposes a new criterion that enables the identification of causal effects not mediated by covariates. In addition, this paper elucidates the mechanism of post-treatment bias, showing that a repeated sequence of nodes can be a potential source of this bias. The results apply to linear data-generating processes represented by directed acyclic graphs with distribution-free error terms.
@@ 1 Introduction
As a generalization of this approach, Pearl1995 developed a framework for causal inference that can also be applied to non-linear causal structures. Within this framework, Pearl1998 applied the back-door criterion—originally proposed by Pearl1993 as a method for covariate selection—to linear data-generating processes represented by directed acyclic graphs (DAGs), showing that it provides a sufficient condition under which a regression coefficient coincides with the total effect of the corresponding explanatory variable on the dependent variable. The back-door criterion has since been generalized by Shpitser2012, VanDerZander2014, Maathuis2015, Perkovic2015, and Perkovic2020. In addition to the back-door criterion, other approaches, such as the front-door criterion (Pearl1995), conditional instrumental variables (BritoandPearl2002), and proxy variables(Kuroki2014; Miao2018) have been established for identifying total effects.
Meanwhile, some previous studies have explored the identification of partial causal relationships. Pearl1998 proposed the single-door criterion for identifying direct effects, and Avin2005 and Shpitser2008 examined the identifiability of path-specific effects. BritoandPearl2002, Tian2009, Foygel2012, Chen2016, and Barber2022 studied the identifiability of path coefficients algebraically by using a system of simultaneous equations based on the graph structure and second-order statistics. Fulcher2020 and Ghassami2025 proposed extensions of the front-door criterion that enable the identification of indirect effects. However, conditions that guarantee a partial regression coefficient coincides with the causal effect not mediated by covariates (i.e., the other explanatory variables) have been identified only in cases where either the direct or total effect is of interest. Consequently, the logical validity of the causal interpretation of partial regression coefficients, as mentioned at the beginning, remains insufficiently established. Given that regression analysis is widely used not only in academic research but also in various practical contexts, it is crucial to provide a theoretically sound basis for its interpretation.
@@
  • Providing a graphical criterion under which a partial regression coefficient represents the causal effect not mediated by the other explanatory variables.
@@
The remainder of this paper is organized as follows. Section 2 introduces the key terms and propositions as preliminaries. Section 3 formulates the problem addressed in this study and presents the main theorem. This section also discusses the mechanisms of post-treatment bias. Section 4 provides illustrative examples. All proofs are presented in Section 5.
@@ 2.4 Controlled Total Effects
Since independence among the error terms is not assumed, this equality cannot be derived based on do-calculus (Pearl2009). However, given that the notation involving the do-operator represents probabilistic concepts of the post-intervention distribution (i.e., the distribution obtained by replacing the right-hand side of the corresponding structural equation with a constant), this interpretation is appropriate. See Example 4.2 in the next section for a concrete illustration of this measure.
@@ 3.2 Main Results
These corollaries were originally established in Spirtes1998 and Pearl1998 based on conditional independence relations and the properties of conditional covariances in linear Gaussian models. However, as also partially noted in Spirtes1998 and Nandy2017, when a constant term is included, the partial regression coefficients in linear structural equation models depend solely on the covariance matrix $(\bm{I}_{N}-\bm{A})^{-1}\mathop{\rm Var}\nolimits(\bm{u})\{(\bm{I}_{N}-\bm{A})^{-1}\}^{\top}$; therefore, their proof does not lose generality. Here, $\bm{I}_{N}$ denotes the $N\times N$ identity matrix. A contribution of the present study is to show that linear causal effects not mediated by covariates can also be consistently estimated using ordinary least squares, even without assuming independence or normality of the error terms in the data-generating process.
@@ 3.3 Post-Treatment Bias
Biases that arise from conditioning on variables affected by a treatment (i.e., descendants in the graph) are referred to as post-treatment bias. Examining the mechanism of this bias helps clarify why condition (i) of the selective-door criterion is required.
@@
As indicated by the back-door criterion, confounding bias can be eliminated by blocking all back-door paths. The first condition (i) of the selective-door criterion implies that $\tau_{pj|do(\bm{S}_{2}\setminus\{X_{j}\})}=0$ and can be understood as a condition to prevent post-treatment bias. Furthermore, this perspective clarifies why merely blocking undirected paths is not sufficient to remove all sources of bias. Since a path is usually defined as a sequence of distinct nodes, the sequence of nodes $X_{j}\rightarrow X_{i}\rightarrow X_{p}\leftarrow X_{i}$ is, for instance, not considered a path. However, the estimated effect of $X_{j}$ on $X_{i}$ can still be distorted through this structure if $X_{p}$ is conditioned on. Another instance of such a structure is provided in Example 4.1 below.